# Connectivity Structure and Dynamics of Nonlinear Recurrent Neural Networks

### v1

@@ VI Single-neuron heterogeneity
This condition is identical to that of an i.i.d. network, but with an average over $\bm{\theta}\sim P(\bm{\theta})$, in addition to the usual Gaussian average over $\eta^{x}(t)$, to account for single-neuron heterogeneity.
@@
To write down the solution the solution for the four-point function, we define new two-frequency correlation functions with an outer average over $\bm{\theta}\sim P(\bm{\theta})$:
@@
along with the usual shorthand $C^{\Phi}_{12}=C^{\Phi}_{\star}(\omega_{1})C^{\Phi}_{\star}(\omega_{2})$. For the unnormalized variables $\Phi_{\bm{\theta}_{i}}(t)$, we obtain
@@
For the normalized variables $\phi_{i}(t)$, we obtain
$\Psi^{\phi}_{\star}=\Bigg{[}|A|^{2}\left(\frac{\mathfrak{C}^{\Phi}_{12}}{C^{\phi}_{12}}+\frac{1}{\alpha\text{PR}^{D}}\frac{C^{\Phi}_{12}}{C^{\phi}_{12}}\right)\\ +\left(A\frac{\mathfrak{C}^{\Phi\phi}_{12}}{C^{\phi}_{12}}+\text{h.c.}\right)+1\Bigg{]}C^{\phi}_{12},$ (92)
where $A={\alpha r_{2}S^{\phi}_{12}}/{(1-\alpha r_{2}\mathfrak{s}^{\phi}_{12})}$.
VI.1 Diverse single-neuron firing rates
An illustrative case of single-neuron heterogeneity occurs when the $\bm{\theta}$ parameters consist of a gain parameter $G_{i}$ for each neuron, modeling diversity in mean firing rates:
$\Phi_{\bm{\theta}_{i}}(x)=G_{i}\phi(x).$ (93)
In this scenario, the dynamics of the network are determined by an effective coupling matrix $\bm{L}\bm{D}\bm{R}^{T}\bm{G}$, where $\bm{G}=\text{diag}(G_{1},\ldots,G_{N})$. This extension of the random-mode model allows us to directly compare the effects of low-dimensional structure in the connectivity (via the component strengths $D_{a}$) with heterogeneity in single-neuron properties (via the gains $G_{i}$). We assume the gains are drawn from a distribution with moments
$q_{n}=\left\langle G^{n}\right\rangle_{G}.$ (94)
To quantify the heterogeneity in firing rates, we define the participation ratio of the gain distribution:
$\text{PR}^{G}=\frac{q_{2}^{2}}{q_{4}}.$ (95)
We analyze collective activity for both the unnormalized activations $\Phi_{\bm{\theta}_{i}}(t)$ and the normalized activations $\phi_{i}(t)$. The latter are analogous to $z$-scored firing rates in neural recordings, where $z$-scoring removes single-neuron heterogeneity while preserving nontrivial structure due to cross-neuron and temporal correlations.
For the unnormalized variables, we calculate the four-point function
$\Psi^{\Phi}_{\star}=\frac{\frac{1}{\text{PR}^{G}}+\frac{1}{\alpha\text{PR}^{D}}|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}{|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}C^{\Phi}_{12},$ (96)
and for the normalized variables,
$\Psi^{\phi}_{\star}=\frac{1+\left(\frac{1}{\text{PR}^{G}}+\frac{1}{\alpha\text{PR}^{D}}-1\right)|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}{|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}C^{\phi}_{12},$ (97)
@@ VI.1 Diverse single-neuron firing rates
$g_{\text{eff}}$ $=\sqrt{\alpha r_{2}q_{2}},$ (98)
$C^{\Phi}_{12}$ $=q_{2}^{2}C^{\phi}_{12}.$ (99)
Eq. 97 demonstrates that reductions in effective rank, described by $\alpha\text{PR}^{D}$, and in the heterogeneity of gains, described by $\text{PR}^{G}$, have symmetric effects on the dimension of normalized activity, with their combined effect captured by the harmonic mean, $\left(\frac{1}{\alpha\text{PR}^{D}}+\frac{1}{\text{PR}^{G}}\right)^{-1}$. This symmetry arises because swapping the distributions of $D_{a}$ and $G_{i}$ does not change the spectrum of $\bm{L}\bm{D}\bm{R}^{T}\bm{G}$ when $\bm{L}$ and $\bm{R}$ are i.i.d., since $\bm{L}\bm{D}\bm{R}^{T}\bm{G}$ and $\bm{R}^{T}\bm{G}\bm{L}\bm{D}$ have the same nonzero eigenvalues. One might expect that normalizing the activations before computing the dimension of activity would remove the effect of heterogeneous gains on the dimension. This is not the case, however, since neuronal gains affect the recurrent dynamics, and not merely the readout.
To further explore this and related effects, we compare the activity dimensions for three sets of activations: (1) unnormalized activations, $\Phi_{\bm{\theta}_{i}}(t)$; (2) normalized activations, $\phi_{i}(t)$; and (3) normalized activations multiplied by an independent set of “readout” gains, $\Phi^{\text{readout}}_{i}(t)=G^{\text{readout}}_{i}\phi_{i}(t)$, where the readout gains have the same distribution as the actual gains. Case (3) corresponds to variables with the same distribution of firing rates as the unnormalized activations, but where this heterogeneity is unrelated to the recurrent dynamics. In this case, we have $C^{\Phi^{\text{readout}}}_{12}=C^{\Phi}_{12}=q_{2}^{2}C^{\phi}_{12}$. The four-point function is
@@
$\Psi^{\Phi^{\text{readout}}}=\Bigg{[}\left(\frac{1}{\text{PR}^{G}}-1\right)\left(|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}+|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}\right)\\ +1+\frac{1}{\alpha\text{PR}^{D}}|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}\Bigg{]}\frac{C^{\Phi^{\text{readout}}}_{12}}{|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}.$ (100)
To validate our theoretical predictions (Eq. 96, Eq. 97, and Eq. 100), we simulated networks with component strengths and gains given by
$D_{a}$ $\propto\exp(-\beta_{D}a/M),$ (101)
$G_{i}$ $\propto\exp(-\beta_{G}i/N).$ (102)
The dimension of activity in these networks is determined by the coupling strength, $g_{\text{eff}}$; the effective rank of the coupling matrix, $\alpha\text{PR}^{D}$; and the participation ratio of the single-neuron gain distribution, $\text{PR}^{G}$. Increases in each of these parameters leads to higher activity dimension for all three sets of activations.
Normalized activations consistently exhibit the highest dimension. While scaling normalized activations by heterogeneous factors reduces dimension as expected, the magnitude of this reduction depends on whether $G_{i}$ or $G^{\text{readout}}_{i}$ is used, even when their distributions are identical. This occurs because neurons with the largest gains preferentially participate in the leading modes of the normalized activations. Further scaling these already-dominant neurons by their gains results in overrepresentation of these neurons compared to using random gains. Thus, the dimension of unnormalized activations is lowest, with the dimension of random-readout activations falling between those of the normalized and unnormalized activations. This effect demonstrates that shuffling the normalizations of recorded neural activities should increase the dimension if firing rates and input covariances are related to one another, as they are in recurrent networks. Not observing such an increase in dimension, on the other hand, would suggest that heterogeneity in firing rates is unrelated to input covariance.
VII Mode overlaps
Figure 10: Input-output mode overlaps and eigenvalue spectra of the fly-brain connectome. (a) Overlap matrix between right and left singular vectors of the connectome. First 1000 modes are shown. (b) First 3000 diagonal elements of the overlap matrix from (a). Horizontal lines indicate an approximate cutoff for $\mathcal{O}(1/\sqrt{N})$ random overlaps. (c) Eigenvalue spectra of various connectome-based matrices. Top: Actual hemibrain spectrum, exhibiting large outliers along the real axis. Middle: Eigenvalues from a realization of the random-mode model, using singular values from the connectome as component strengths $D_{a}$, resulting in a circularly symmetric spectrum. Bottom: Random-mode model realization with input-output mode overlaps incorporated, where overlaps are given by the inner products of left and right singular vectors from the fly-connectome SVD. Data from Scheffer et al. [21], $N=18028$ neurons.
Thus far, we have assumed that the matrices $\bm{L}$ and $\bm{R}$ are independent of each other. Under this assumption, the overlap $\bm{L}_{a}^{T}\bm{R}_{b}$ between input mode $b$ and output mode $a$ is a random, $\mathcal{O}(1/\sqrt{N})$ quantity. However, to implement computations, modes must interact, suggesting that the coupling matrices of real neural circuits have larger, structured overlaps between input and output modes.
To examine interactions between different modes in a real neural circuit, we returned to the connectome of Drosophila melanogaster. We computed the overlap matrix between its right and left singular vectors, which correspond to the input and output modes, respectively. Rather than the unstructured, $\mathcal{O}(1/\sqrt{N})$ overlaps expected from i.i.d. random modes, we observe structured, large overlaps (Fig. 10a). These large overlaps are particularly pronounced near the diagonal of the overlap matrix (Fig. 10b).
Motivated by these observations, we now extend the random-mode model to incorporate structured overlaps between input and output modes. Specifically, we define $\bm{L}$ and $\bm{R}$ as zero-mean Gaussian matrices with the following covariance structure:
$\left\langle L_{ia}L_{jb}\right\rangle$ $=\delta_{ij}\delta_{ab}\frac{1}{N},$ (103)
$\left\langle R_{ia}R_{jb}\right\rangle$ $=\delta_{ij}\delta_{ab}\frac{1}{N},$ (104)
$\left\langle L_{ia}R_{jb}\right\rangle$ $=\delta_{ij}\delta_{ab}\frac{O_{ab}}{N},$ (105)
where $O_{ab}$ is the overlap matrix. In this case,
$\left\langle\bm{L}^{T}\bm{R}\right\rangle=\bm{O}.$ (106)
To capture the most salient feature of the fly connectome, large diagonal overlaps, we specialize to a diagonal form for the overlap matrix:
$O_{ab}=\delta_{ab}\rho_{a},$ (107)
where $\rho_{a}$ is the correlation between the $a$-th input and output modes. This form of the overlaps implies that, for large $N$, $\bm{L}_{a}^{T}\bm{R}_{a}=\rho_{a}$, while $\bm{L}_{a}^{T}\bm{R}_{b}$ for $a\neq b$ remains a random $\mathcal{O}(1/\sqrt{N})$ quantity. Incorporating this structure allows the random-mode model to more accurately match the eigenspectrum of the connectome, which has large outliers along the real axis (Fig. 10c).
Defining an instance of this connectivity with a well-defined large-$N$ limit requires specifying a joint distribution from which $(D_{a},\rho_{a})$ are sampled for each $a$. Our analytic expressions for the two- and four-point functions depend on averages $\left\langle\cdots\right\rangle_{D,\rho}$ over this distribution. Note that for large $N$, self-connections (autapses) in $\bm{J}$ are given by
$J_{ii}=\alpha\left\langle D\rho\right\rangle_{D,\rho},$ (108)
which is, in general, $\mathcal{O}(1)$. To avoid biologically unrealistic strong self-connections, one can choose $\left\langle D\rho\right\rangle_{D,\rho}=0$, in which case $J_{ii}$ is random and $\mathcal{O}(1/\sqrt{N})$. Alternatively, if one interprets each model neuron as a cluster of biological neurons, strong self-connections can be interpreted as aggregated within-cluster connectivity [64, 65]. Furthermore, note that the covariance between reciprocal pairs of couplings is given by $N\left\langle J_{ij}J_{ji}\right\rangle_{\bm{J}}=\sigma$ ($i\neq j$) where
$\sigma=\alpha\left\langle(D\rho)^{2}\right\rangle_{D,\rho}.$ (109)
We compute the two- and four-point functions using the path integral (Appendix I). As in previous cases, we obtain single-site processes for both neurons and latent variables. The neuronal single-site process is given by
$T[x](t)=\eta^{x}(t)+[R_{\star}\ast\phi](t),$ (110)
where $\eta^{x}\sim\mathcal{GP}(0,Q_{\star})$ and $\ast$ denotes convolution. The input covariance $Q_{\star}(t_{1},t_{2})$ and convolutional kernel $R_{\star}(t_{1},t_{2})$ are defined in terms of statistics of the latent-variable single-site process. Due to the linearity of the latent-variable dynamics, we can express them as
We enforce $C^{\phi}_{\star}(t_{1},t_{2})=\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{\star}$ and $S^{\phi}_{\star}(t_{1},t_{2})=\left\langle{\delta\phi(t_{1})}/{\delta I(t_{2})}\right\rangle_{\star}$, yielding a self-consistent system that determines $C^{\phi}_{\star}(t_{1},t_{2})$ and $S^{\phi}_{\star}(t_{1},t_{2})$. Unlike previous cases, there is no effective coupling strength $g_{\text{eff}}$ describing the relationship between $Q_{\star}(t_{1},t_{2})$ and $C^{\phi}_{\star}(t_{1},t_{2})$; instead, this relationship involves a convolution with a kernel defined using the joint distribution over $(D,\rho)$. This can be viewed as a frequency-dependent effective coupling strength.
Expanding the self-coupling kernel $R_{\star}(\omega)$ (Eq. 112) in powers of $\rho$ yields the first two terms
$R_{\star}(\omega)=J_{ii}+\sigma S^{\phi}_{\star}(\omega)+\cdots$ (113)
The first term is what we get in the single-site picture for i.i.d. cross-neuron couplings and deterministic $\mathcal{O}(1)$ self couplings. The second term appears in the single-site picture if there are correlations between reciprocal pairs of couplings, but the structure is otherwise i.i.d. (Appendix E). Nonzero higher-order terms reflect that the structure of the random-mode model is not fully captured by either of these forms. Note that the first two terms being zero forces all subsequent terms to be zero. Thus, nonzero covariance among reciprocal pairs of couplings is a necessary and sufficient condition for attaining a nonzero self-coupling kernel in the single-site picture.
Figure 11: Effect of input-output mode correlations on dimension of activity in the random-mode model. Dimension of activity $\text{PR}^{\phi}$ versus correlation parameter $\rho_{1}$ in the $(D,\rho)$ joint distribution, for various coupling strengths $g_{\text{eff}}$ and effective ranks $\alpha\text{PR}^{D}$. Thin dots: individual simulations; thick dots: means over 10 simulations; lines: theoretical predictions. All simulations use $N=5000$ neurons.
To compute the four-point function, we write down and invert the frequency-dependent Hessian, and obtain
$\Psi^{\phi}_{\star}$ $=\frac{V-1+\left\{\frac{(1+X_{12})(1+X_{21}^{*})}{1-W^{*}}+\text{h.c.}\right\}}{|1-U|^{2}}C^{\phi}_{12},$ (114a)
$U$ $=\alpha\left\langle D^{2}\Sigma_{1}\Sigma_{2}\right\rangle_{D,\rho},$ (114b)
$V$ $=\alpha\left\langle D^{4}|\Sigma_{1}\Sigma_{2}|^{2}\right\rangle_{D,\rho},$ (114c)
$W$ $=\alpha\left\langle D^{2}\rho^{2}\Sigma_{1}^{*}\Sigma_{2}\right\rangle_{D,\rho},$ (114d)
$X_{12}$ $=\alpha\left\langle D^{3}\rho\Sigma_{1}|\Sigma_{2}|^{2}\right\rangle_{D,\rho},$ (114e)
$\Sigma_{k}$ $=\frac{S^{\phi}_{k}}{1-D\rho S^{\phi}_{k}}\>\>\text{for}\>\>k=1,2.$ (114f)
We validated this theory by choosing parameterizations of $D_{a}$ and $\rho_{a}$. We used $\alpha=M/N=1$ and $D_{a}\propto\exp(-\beta_{D}a/N)$ for the component strengths, as before. For the overlaps $\rho_{a}$, we use a sigmoidal shape inspired by Fig. 10b in which the leading components have larger input-output mode overlaps. In particular, we used
$\rho_{a}=\rho_{1}\left(1-\frac{2}{1+\exp(-\beta_{\rho}(a/N-u_{\text{flip}}))}\right),$ (115)
which varies from $\rho_{1}$ to $-\rho_{1}$, hitting $\rho_{a}=0$ at $a/N=u_{\text{flip}}$. We fixed $\beta_{\rho}=16$, varied $\rho_{1}$, and set $u_{\text{flip}}$ such that $\left\langle D\rho\right\rangle_{D,\rho}=0$, eliminating strong self couplings. Across values of $g_{\text{eff}}$ and $\text{PR}^{D}$, sufficiently strong overlaps of either sign reduce activity dimension (Fig. 11). The maximum dimension occurs for $\rho_{1}<0$, a regime in which activity projected along connectivity modes with the greatest strength is most strongly suppressed by the recurrent dynamics. This leftward shift is reduced for larger $g_{\text{eff}}$.
VIII Discussion
A key question in neuroscience and machine learning is how learning shapes collective dynamics through modifications of a neural network’s connectivity. While our analysis has focused on disordered networks that generate chaotic activity and are not trained to perform tasks, the smooth low-rank connectivity structure we have explored provides a starting point for understanding the impact of learned coupling matrices on collective dynamics. Recent works have shown that large neural networks trained on real-world tasks exhibit smooth spectral structures in their coupling matrices, similar to those we have studied [20]. Moreover, while networks trained on simple tasks often display low-rank updates to their initial couplings, with a finite number of outlier eigenvalues emerging from the bulk [19, 46], training networks on many tasks, as is increasingly common, may produce coupling matrices with a low but extensive number of significant components [66, 67]. In such scenarios, the random-mode model, or extensions of it, may predict the resulting activity dimension and other aspects of collective dynamics. The path-integral formalism we have developed is also amenable to analyzing cases where the coupling matrix is drawn from a Gibbs ensemble with an energy function equal to the loss function for a task. This approach represents scenarios where couplings are trained to minimize a loss function with stochastic gradient descent [52]. Finally, initializing couplings with low-rank structure could improve training procedures like FORCE learning, which are based on suppressing chaos, because the initial strange attractor is lower-dimensional [11, 12].
Historically, neural connectivity was often characterized through network motifs – recurring patterns like cycles or chains among small neuronal groups [68, 14, 69, 70]. Such local statistics can be estimated using connectivity datasets obtained from methods that sample subsets of synapses. However, recent advances in whole-brain connectome reconstruction have unveiled global synaptic connectivity structures [71]. These connectomes reveal a more complex picture, where the spectrum and modes of the connectivity matrix exhibit large-scale structure that cannot be explained by local motifs [72]. Our parameterization of coupling matrices attempts to describe this global perspective, offering control over both the spectrum and mode overlaps. This could be particularly valuable in interpreting large-scale recordings of neural activity in the context of known anatomical connectivity.
In a similar vein, Tiberi et al. [72] proposed a random matrix theory approach for noise-driven linear networks based on parameterizing the eigendecomposition of $\bm{J}$. By contrast, our use of a DMFT approach allows us to describe nonlinear dynamics. Their parameterization also does not capture relationships between eigenvalues and eigenvectors, which are likely crucial for computations. Our parameterization, which is more akin to an SVD than an eigendecomposition, allows for a simultaneous parameterization of component strengths and corresponding mode overlaps. Other parameterizations of connectivity similar to the random-mode model have been used to study the relationship between spiking and rate-based recurrent networks [73] and to train spiking networks more efficiently [74].
Our description of single-neuron heterogeneity, where each neuron is parameterized by an arbitrary vector $\bm{\theta}$, can be readily applied to explore how various forms of heterogeneity influence dimensionality and other features of collective activity. Introducing specific forms of heterogeneity constrained by observed electrophysiological properties [75] and other cell-specific characteristics across neurons could provide insight on the functional consequences of neuronal diversity [76].
On the technical side, we have demonstrated that the two-site cavity method and the analysis of fluctuations around the saddle point in the path integral approach yield equivalent results for computing the four-point function. This equivalence implies that the intricate cavity construction is equivalent to computing and inverting the Hessian of the action in the path integral. A deeper understanding of the relationship between these methods could lead to new hybrid approaches that combine the strengths of both techniques.
In conclusion, the random-mode model and associated analytical techniques provide a set of tools for exploring the relationship between connectivity and collective dynamics. Our parameterization of connectivity is a middle ground between deterministic low-rank models and completely random networks. Future work incorporating more complex mode interactions, diverse neuronal properties, and learning dynamics will further bridge neural connectivity, activity, and function.
Acknowledgements
We gratefully acknowledge L.F. Abbott for discussions and feedback on this work.
Appendix A Glossary of Terms
Number of neurons ($N$)
Total number of neurons in the network.
Number of modes ($M$)
Number of components in the random-mode model.
(Effective) coupling strength ($g$, $g_{\text{eff}}$)
Standard deviation of the couplings times $\sqrt{N}$ in the i.i.d. connectivity model (Eq. 14), random-mode model (Eq. 74), or variants thereof (Eq. 98).
Random-mode model
Network connectivity model: $\bm{J}=\sum_{a=1}^{M}D_{a}\bm{L}_{a}\bm{R}_{a}^{T}$ (Eq. 22).
Connectivity component ($\bm{L}_{a}\bm{R}_{a}^{T}$)
Outer product of output and input modes.
Input mode ($\bm{R}_{a}$)
Output mode ($\bm{L}_{a}$)
Component strength ($D_{a}$)
Scaling factor for the $a$-th component.
Ratio of modes to neurons ($\alpha$)
Given by $\alpha=M/N$ (Eq. 24).
Participation ratio of component strengths ($\text{PR}^{D}$)
Effective rank ($\alpha\text{PR}^{D}$)
Connectivity dimensionality measure (Eq. 27).
Two-point function
Correlation function given by $C^{\phi}(t_{1},t_{2})=\frac{1}{N}\sum_{i}\langle\phi_{i}(t_{1})\phi_{i}(t_{2})\rangle_{t}$ (Eq. 5).
Four-point function
Higher-order correlation function given by $\Psi^{\phi}(\tau_{1},\tau_{2})=\frac{1}{N}\sum_{i,j}C^{\phi}_{ij}(\tau_{1})C^{\phi}_{ij}(\tau_{2})$ (Eq. 7). See also time-by-time definition (Eq. 9).
Appendix B Preprocessing of the fly hemibrain connectome
We used the dataset originally presented in Scheffer et al. [21]. To form the connection matrix for analysis, we used synaptic counts as a measure of connection strength, determining the magnitudes of the matrix elements. Neurons were assigned neurotransmitter probabilities based on a machine-learning analysis applied to the electron-microscopy data [22]. We selected the most probable neurotransmitter assignment for each neuron. The neurotransmitter types were then mapped to signs (excitatory or inhibitory) according to Table 1.
To normalize the connectivity matrix, we used an iterative scaling procedure analogous to the Sinkhorn–Knopp algorithm that proceeds as follows. We first rescale the positive elements in each row by a row-specific factor, such that the L2-norm of the positive elements is $1/\sqrt{2}$. The same procedure is applied to the negative elements, using a separate row-specific factor. This results in each row having a L2-norm of one with balanced excitatory and inhibitory inputs. A small number of neurons with exclusively excitatory or inhibitory inputs are left with an L2-norm of $1/\sqrt{2}$ instead of one. Following the row normalization, we perform the same operations on the columns. This disrupts the previous row normalization. Thus, we iterate the row and column normalization steps in an alternating fashion until convergence was achieved. The final result is a matrix with proper scaling along both rows and columns, which we interpret as the projection of the original matrix onto the set of properly row- and column-normalized matrices. Importantly, it has the same nonzero elements with the same signs as the original matrix.
Neurotransmitter Synaptic Sign
GABA -1 (Inhibitory)
Acetylcholine +1 (Excitatory)
Glutamate -1 (Inhibitory)
Serotonin Ignored
Octopamine Ignored
Dopamine Ignored
Neither Ignored
Table 1: Mapping of neurotransmitters to synaptic signs
Appendix C Temporal limits and inverses
We start by examining the properties of $U(t_{1},t_{2},t_{3},t_{4})$. $U$ vanishes if one of the four time points is far (measured on the timescale of $C^{\phi}_{\star}(\tau)$) from the other three. For $t_{1}$ or $t_{2}$, this is a consequence of the perturbation associated with $\hat{x}(t_{1})$ or $\hat{x}(t_{2})$ being far away; for $t_{3}$ or $t_{4}$, this is a consequence of $\langle\phi(t_{i})\rangle_{\star}=0$ separating from the expectation. $U$ also vanishes if there is more than one time point which is far from the remaining ones by the same arguments. Thus, the time points need to be close in pairs. However, if $t_{1},t_{2}$ are close and $t_{3},t_{4}$ are close, but the pairs are distant, $U$ still vanishes since the perturbations associated with $\hat{x}(t_{i})$ are both far. This leaves three possibilities for a non-vanishing $U$: $t_{1},t_{3}$ close and $t_{2},t_{4}$ close, but the pairs are distant; $t_{1},t_{4}$ close and $t_{2},t_{3}$ close, but the pairs are distant; or all time points are close.
This justifies taking the limits Eq. 34 for $K$.
Appendix D Incorporating sources
This section outlines the general approach for incorporating source terms and then applies it to the specific case of correlations among the pre-activations. Our goal is to compute
$Z=\int\mathcal{D}\bm{C}\int\mathcal{D}U\int\mathcal{D}\hat{U}\exp(-N\tilde{\mathcal{S}}[\bm{C},U,\hat{U}]),$ (123)
@@ Appendix D Incorporating sources
$\tilde{\mathcal{S}}[\bm{C},U,\hat{U}]=\hat{U}U+\mathcal{S}[\bm{C};U].$ (124)
where
The fluctuations of $\hat{U}$ are given by the bottom-right element of the inverse Hessian evaluated at the saddle point. Using a Schur complement to compute this element, we obtain the expression Eq. 54 stated in the main text.
$\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}=-C_{12}^{x},\quad\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}=\begin{pmatrix}g^{2}S_{12}^{x}\\ -C_{12}^{x\phi}\end{pmatrix},$ (128)
Appendix E Partially symmetric disorder
$\left\langle J_{ij}J_{k\ell}\right\rangle_{\bm{J}}=g^{2}\delta_{ik}\delta_{j\ell}+g^{2}\rho\delta_{i\ell}\delta_{jk}.$ (130)
$Z[\bm{X},\bm{Y}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\exp\bigg{(}i\sum_{i}\hat{x}_{i}T[x_{i}]-i\sqrt{{g^{2}-|\sigma|}}\sum_{ij}X_{ij}\hat{x}_{i}\phi_{j}-i\sqrt{\frac{|\sigma|}{2}}\sum_{ij}Y_{ij}(\hat{x}_{i}\phi_{j}+\text{sgn}(\sigma)\phi_{i}\hat{x}_{j})\bigg{)}.$ (132)
@@ Appendix E Partially symmetric disorder
We obtain, for the disorder-averaged path integral,
$\hat{C}^{\phi}_{\star}(t_{1},t_{2})$ $=0,$ (142)

### v2

@@ I.2 Four-point functions
The self-consistency condition Eq. 155c is identical to that of an i.i.d. network, but with an average over $\bm{\theta}\sim P(\bm{\theta})$, in addition to the usual Gaussian average over $\eta^{x}(t)$, to account for single-neuron heterogeneity.
J.2 Four-point functions
@@
To write down the solution for the four-point function, we define new two-frequency correlation functions with an outer average over $\bm{\theta}\sim P(\bm{\theta})$:
@@
along with the usual shorthand $C^{\Phi}_{12}=C^{\Phi}_{\star}(\omega_{1})C^{\Phi}_{\star}(\omega_{2})$.
The frequency-dependent Hessian at the saddle point is
$\bm{H}=\left(\begin{array}[]{cc|cc}0&\alpha r_{2}&0&-1\\ \alpha r_{2}&-\alpha r_{4}{C_{{12}}^{\Phi}}&-1&0\\ \hline\cr 0&-1&0&\mathfrak{s}^{\Phi}_{12}\\ -1&0&(\mathfrak{s}^{\Phi}_{12})^{*}&-\mathfrak{C}^{\Phi}_{12}\end{array}\right),$ (157)
where the new $\mathfrak{F}$raktur variables are defined in Eq. 156. We invert the Hessian using Eq. 146 and identify the upper-left element as $\Phi_{\bm{\theta}_{i}}(\bm{\omega})$:
@@
Combining the inverted Hessian with the relevant quantities for computing fluctuations of normalized variables,
$\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}=-C^{\phi}_{12},\>\>\>\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}=\begin{pmatrix}0\\ 0\\ (S^{\phi}_{12})^{*}\\ -\mathfrak{C}^{\Phi\phi}_{12}\end{pmatrix}$, (159)
@@
we obtain for the normalized variables $\phi_{i}(t)$:
@@
$\Psi^{\phi}_{\star}=\Bigg{[}|A|^{2}\left(\frac{\mathfrak{C}^{\Phi}_{12}}{C^{\phi}_{12}}+\frac{1}{\alpha\text{PR}^{D}}\frac{C^{\Phi}_{12}}{C^{\phi}_{12}}\right)+\left(A\frac{\mathfrak{C}^{\Phi\phi}_{12}}{C^{\phi}_{12}}+\text{h.c.}\right)+1\Bigg{]}C^{\phi}_{12},$ (162)
@@
$A=\frac{\alpha r_{2}S^{\phi}_{12}}{1-\alpha r_{2}\mathfrak{s}^{\phi}_{12}}.$ (163)
J.3 Four-point functions for firing-rate heterogeneity
Upon specializing the above general results to the case where $\bm{\theta}_{i}$ contains a single gain parameter, $G_{i}$, we obtain the four-point functions for unnormalized (Eq. 46a) and normalized (Eq. 46b) variables given in the main text.
Note that, like the distribution over component strengths $D_{a}$, two-point functions of activity depend only on the second moment of $G_{i}$, and four-point functions depend only on the second and fourth moments of $G_{i}$.
For “random readout” gains with the same distribution as the recurrent gains (see main text), the four-point function is
$\Psi^{\Phi^{\text{readout}}}=\Bigg{[}\left(\frac{1}{\text{PR}^{G}}-1\right)\left(|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}+|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}\right)+1+\frac{1}{\alpha\text{PR}^{D}}|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}\Bigg{]}\frac{C^{\Phi^{\text{readout}}}_{12}}{|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}},$ (164)
where $C^{\Phi^{\text{readout}}}_{12}=C^{\Phi}_{12}=q_{2}^{2}C^{\phi}_{12}$
We conclude with a note on gain modulation in recurrent neural networks. It has been shown that modulating single-neuron gains while keeping synaptic weights fixed has great expressive power over the dynamics of recurrent neural networks [95]. Here, we consider random gains rather than judiciously chosen or learned gains. In particular, they are sampled independently of the connectivity. However, our framework could be extended to model gains that are related to the connectivity and, more ambitiously, could be extended to model learned gain parameters (or, as explored in the Discussion, learned connectivity).
Appendix K Path-integral calculation of two- and four-point functions for the random-mode model with diagonally structured overlaps
To generate matrices $\bm{L}$ and $\bm{R}$ with appropriate correlation structure, we express them in terms of independent Gaussian random matrices $\bm{X}^{1},\bm{X}^{2},$ and $\bm{Y}$, each with variance $1/N$:
${L}_{ia}$ $=\sqrt{{1-|\rho_{a}|}}{X}^{1}_{ia}+\sqrt{{|\rho_{a}|}}{Y}_{ia},$ (165)
$C^{\phi}(t_{1},t_{2})$ $=\frac{1}{N}\sum_{i=1}^{N}\phi_{i}(t_{1})\phi_{i}(t_{2}),$ $\hskip-72.26999pt\text{conjugate: }\hat{Q}(t_{1},t_{2});$ (167)
$S^{\phi}(t_{1},t_{2})$ $=\frac{1}{N}\sum_{i=1}^{N}\frac{\delta\phi_{i}(t_{1})}{\delta I_{i}(t_{2})},$ $\hskip-72.26999pt\text{conjugate: }\hat{R}(t_{1},t_{2});$ (168)
$Q(t_{1},t_{2})$ $=\frac{1}{N}\sum_{a=1}^{M}z_{a}(t_{1})z_{a}(t_{2}),$ $\hskip-72.26999pt\text{conjugate: }\hat{C}^{\phi}(t_{1},t_{2});$ (169)
$R(t_{1},t_{2})$ $=\frac{1}{N}\sum_{a=1}^{M}D_{a}\rho_{a}\frac{\delta z_{a}(t_{1})}{\delta I_{a}(t_{2})},$ $\hskip-72.26999pt\text{conjugate: }\hat{S}^{\phi}(t_{1},t_{2}).$ (170)
[…]
@@
K.1 Two-point functions
The saddle-point conditions yield:
$\hat{C}^{\phi}(t_{1},t_{2})$ $=\hat{Q}_{\star}(t_{1},t_{2})=0,$ (184)